# A Practical Introduction to Kernel Discrepancies: MMD, HSIC & KSD

### v2

@@ 1 Kernels
Given two Hilbert spaces ${\mathcal{H}}_{1}$ and ${\mathcal{H}}_{2}$ with associated inner products $\langle\cdot,\cdot\rangle_{{\mathcal{H}}_{1}}$ and $\langle\cdot,\cdot\rangle_{{\mathcal{H}}_{2}}$, respectively, the tensor product $\otimes$ and Hilbert–Schmidt inner product $\langle\cdot,\cdot\rangle_{\mathrm{HS}}$ are defined as $\langle f_{1}\otimes f_{2},g_{1}\otimes g_{2}\rangle_{\mathrm{HS}}=\langle f_{1},g_{!}\rangle_{{\mathcal{H}}_{1}}\langle g_{1},g_{2}\rangle_{{\mathcal{H}}_{2}}$ for all $f_{1},g_{1}\in{\mathcal{H}}_{1}$ and all $f_{2},g_{2}\in{\mathcal{H}}_{2}$. Let $P_{XY}$ be a joint probability distribution on ${\mathcal{X}}\times{\mathcal{Y}}$, and consider kernels $k^{\mathcal{X}}$} and $k^{\mathcal{Y}}$} on ${\mathcal{X}}\times{\mathcal{X}}$ and ${\mathcal{Y}}\times{\mathcal{Y}}$, respectively. The cross-covariance operator is defined as the linear operator ${\mathcal{C}}_{P_{XY}}\colon{\mathcal{H}}_{{k^{\mathcal{Y}}}}\to{\mathcal{H}}_{{k^{\mathcal{X}}}}$ satisfying
@@ 2.3 KSD: Kernel Stein Discrepancy
$\mathrm{KSD}_{{\mathcal{A}}_{P}^{\prime}}(P,Q)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sup_{{\bm{f}}\in{\mathcal{F}}}\ \mathbb{E}_{Q}\!\left[({\mathcal{A}}_{P}^{\prime}{\bm{f}})(X)\right]\leavevmode\nobreak\ =\leavevmode\nobreak\ \sup_{f_{1}\in{\mathcal{H}}\,:\,\|f_{1}\|_{{\mathcal{H}}}\leq 1}\ \mathbb{E}_{Q}\!\left[f_{1}\right]-\mathbb{E}_{P}\!\left[f_{1}\right]\leavevmode\nobreak\ =\leavevmode\nobreak\ \mathrm{MMD}_{k}(P,Q),$ (59)
@@ 4 Kernel pooling: adaptive kernel discrepancies estimators
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\leavevmode\nobreak\ S_{k}/\sigma_{k}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\ S_{k}/\sigma_{k}.$ (74)
@@
$\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\leavevmode\nobreak\ S_{k}/\sigma_{k}\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{1}{\nu}\log\left(\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\exp\left(\nu S_{k}/\sigma_{k}\right)\right)$ (77)
@@
$\max_{k\in{\mathcal{K}}}S_{k}/\sigma_{k}-\frac{\log(|{\mathcal{K}}|)}{\nu}\leavevmode\nobreak\ \leavevmode\nobreak\ \leq\leavevmode\nobreak\ \leavevmode\nobreak\ \underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\leavevmode\nobreak\ S_{k}/\sigma_{k}\leavevmode\nobreak\ \leavevmode\nobreak\ \leq\leavevmode\nobreak\ \leavevmode\nobreak\ \max_{k\in{\mathcal{K}}}S_{k}/\sigma_{k}$ (78)

### v3

@@ 1 Kernels
Given two Hilbert spaces ${\mathcal{H}}_{1}$ and ${\mathcal{H}}_{2}$ with associated inner products $\langle\cdot,\cdot\rangle_{{\mathcal{H}}_{1}}$ and $\langle\cdot,\cdot\rangle_{{\mathcal{H}}_{2}}$, respectively, the tensor product $\otimes$ and Hilbert–Schmidt inner product $\langle\cdot,\cdot\rangle_{\mathrm{HS}}$ are defined as $\langle f\otimes g,\tilde{f}\otimes\tilde{g}\rangle_{\mathrm{HS}}=\langle f,\tilde{f}\rangle_{{\mathcal{H}}_{1}}\langle g,\tilde{g}\rangle_{{\mathcal{H}}_{2}}$ for all $f,\tilde{f}\in{\mathcal{H}}_{1}$ and all $g,\tilde{g}\in{\mathcal{H}}_{2}$. Let $P_{XY}$ be a joint probability distribution on ${\mathcal{X}}\times{\mathcal{Y}}$, and consider kernels $k^{\mathcal{X}}$} and $k^{\mathcal{Y}}$} on ${\mathcal{X}}\times{\mathcal{X}}$ and ${\mathcal{Y}}\times{\mathcal{Y}}$, respectively. The cross-covariance operator is defined as the linear operator ${\mathcal{C}}_{P_{XY}}\colon{\mathcal{H}}_{{k^{\mathcal{Y}}}}\to{\mathcal{H}}_{{k^{\mathcal{X}}}}$ satisfying
@@ 2.3 KSD: Kernel Stein Discrepancy
$\mathrm{KSD}_{{\mathcal{A}}_{P}^{\prime}}(P,Q)\penalty 10000\ =\penalty 10000\ \sup_{{\bm{f}}\in{\mathcal{F}}}\ \mathbb{E}_{Q}\!\left[({\mathcal{A}}_{P}^{\prime}{\bm{f}})(X)\right]\penalty 10000\ =\penalty 10000\ \sup_{f_{1}\in{\mathcal{H}}\,:\,\|f_{1}\|_{{\mathcal{H}}}\leq 1}\ \mathbb{E}_{Q}\!\left[f_{1}\right]-\mathbb{E}_{P}\!\left[f_{1}\right]\penalty 10000\ =\penalty 10000\ \mathrm{MMD}_{k}(P,Q),$ (59)
@@ 4 Kernel pooling: adaptive kernel discrepancies estimators
$\underset{k\in{\mathcal{K}}}{\mathrm{mean}}\penalty 10000\ S_{k}/\sigma_{k}\penalty 10000\ =\penalty 10000\ \frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\ S_{k}/\sigma_{k}.$ (74)
@@
$\underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\penalty 10000\ S_{k}/\sigma_{k}\penalty 10000\ =\penalty 10000\ \frac{1}{\nu}\log\left(\frac{1}{|{\mathcal{K}}|}\sum_{k\in{\mathcal{K}}}\exp\left(\nu S_{k}/\sigma_{k}\right)\right)$ (77)
@@
$\max_{k\in{\mathcal{K}}}S_{k}/\sigma_{k}-\frac{\log(|{\mathcal{K}}|)}{\nu}\penalty 10000\ \penalty 10000\ \leq\penalty 10000\ \penalty 10000\ \underset{k\in{\mathcal{K}}}{\mathrm{fuse}}\penalty 10000\ S_{k}/\sigma_{k}\penalty 10000\ \penalty 10000\ \leq\penalty 10000\ \penalty 10000\ \max_{k\in{\mathcal{K}}}S_{k}/\sigma_{k}$ (78)